OpenStudy (anonymous):

Determine the geometric mean of 5 and 8.

OpenStudy (anonymous):

\[\frac{ 5 }{ x } = \frac{ x }{ 8 }\]

OpenStudy (anonymous):

40=4*10 sqrt40=sqrt(4*10) sqrt40=sqrt4*sqrt10 sqrt40=2*sqrt10

OpenStudy (anonymous):

Cross multiply. \[x^{2} = 40\] Where does the 10 come in? Did you divide?

OpenStudy (anonymous):

2 sqrt(10)~~6.32456

OpenStudy (anonymous):

I see. You eliminated the exponent and then square root 40. Okay, got it. :)